A black hole has a Schwarzschild radius of 88.61 km or 8.86×104 meters. What is its mass?
Answer #1

By using the mathematical formula

RS= 2GMBH/c2

Mass of the black holeMBH = (RS X c2)/ 2G

Where RS is the Schwarzschild radius

MBH= mass of Black hole

c= velocity of light

G= gravitational constant= 6.674×1011 m3⋅kg−1⋅s−2.

Value of c in m/s= 186,000 mi/s

MBH= (8.86×104 X 186,000 X 186,000)/2X 6.674×1011

=30.567X 1014/13.348 X 1011

=2.29 X 1025 kg will be the mass of the Blackhole